### v1

# Superconducting Diode Effect in Weak Localization Regime

Abstract
We study a dirty two-dimensional superconductor with Rashba spin-orbit coupling and in-plane Zeeman fields described by the nonlinear sigma model that includes short-range electron-electron interactions from the Coulomb and Cooper channels. The renormalized Ginzburg-Landau theory, which includes the weak localization effects at the one-loop level, is constructed by using the Keldysh functional formalism. It is shown that the tricritical point appears in the phase diagram. The superconducting diode quality factor increases divergently as the system approaches the tricritical point. Near the superconducting phase transition lines, the absolute value of the diode quality factor decreases due to the cooperation of localization and interactions. The normal conductivity of the resistive state, in which the superconducting state is destroyed by the critical current, is calculated, and localization behaviors are demonstrated.
Introduction: The phenomenon in which the critical currents of superconductors become nonreciprocal, known as the superconducting diode (SD) effect [1, 2], has attracted much attention. Theoretical studies have been carried out for various types of noncentrosymmetric superconductors [3, 4, 5, 6, 7, 8, 9]. Not only clean systems but also systems with uncorrelated impurities have been studied in detail at the mean-field level [10]. If electron systems are disordered enough, they are well described by the nonlinear sigma model (NLSM) [11, 12, 13, 14], and dirty superconductors can also be described by the NLSM [15, 16, 17, 18]. Recently, an NLSM for a dirty superconductor with spin-orbit coupling (SOC) and a Zeeman field was derived [19] and used to investigate the SD effect in Josephson junctions at the mean-field level [20]. However, the Anderson localization effect, a correction to the mean-field theory in dirty systems, has not yet been highlighted in studies of the SD effect.
The role of Anderson localization in superconductors is an important topic of ongoing debate. Without considering strong mesoscopic correlations and electron-electron (e-e) interactions, the properties of dirty $s$-wave superconductors at zero magnetic field remain unchanged with respect to the strength of non-magnetic disorder in the weak localization (WL) regime, known as the Anderson theorem [21]. If the Coulomb interaction is concerned, the transition temperature for the $s$-wave pairing states is suppressed in the WL regime [22, 23, 24]. The opposite effect can occur in the strong localization regime by the multifractality effect [25, 26, 27, 28]. It was also shown that, in the WL regime, the Coulomb interaction suppresses the critical current of conventional dirty superconductors at zero magnetic field and low temperatures [29, 30].
In this Letter, we study the interplay of the e-e interaction and disorder on the SD effect and clarify the WL correction to conductivity when the superconducting (SC) state is broken by the critical current. We work with the NLSM for a dirty two-dimensional (2D) noncentrosymmetric superconductor, including the interactions from the Coulomb and Cooper channels in the short-range limit, and establish a renormalized Ginzburg-Landau (GL) theory at the one-loop level. While many studies [28, 31, 32, 33] have been conducted in the replica formalism, we formulate the theory with the Keldysh functional formalism [14, 15, 17, 34].
Model: We consider a dirty 2D superconductor aligned in the $xy$-plane with the Rashba SOC and a Zeeman field in the $x$-direction. In the Keldysh functional formalism, the partition function is given as $Z=\int D[Q,\Delta]\exp\left({iS[Q,\Delta]}\right)$ with [19]
$S[\Delta,Q]$ $=S_{\rm HS}[\Delta]+S_{\rm N}[\Delta,Q]+S_{\rm T}[Q],$ (1)
where $S_{\rm HS}$ is responsible for the Hubbard-Stratonovich transformation associated with the SC state, $S_{\rm N}$ is the conventional non-interacting part, and $S_{\rm T}$ is the topological part. Their explicit forms are shown in Supplemental material [35]. A $\times 8$ 8 matrix field $Q$ satisfies the conditions $Q^{2}=1$ and $\mathrm{Tr}\,\{Q\}=0$. Let $\gamma_{\mu},~\tau_{\mu}$, and $\sigma_{\mu}~(\mu=0,\dots,4)$ be the Pauli matrices in the Keldysh, Nambu, and spin spaces, respectively. In particular, we write $\gamma_{\rm c}=\gamma_{0}$ and $\gamma_{\rm q}=\gamma_{1}$ and often abbreviate $\gamma_{0}$ for simplicity. The order parameter of superconductivity, that is, the SC gap, is then described by the matrix $\gamma_{\rm c}\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q}$, where $\check{\Delta}_{\rm c(q)}=\tau_{+}\Delta_{\rm c(q)}+\tau_{-}\Delta^{*}_{\rm c(q)}$ with $\tau_{\pm}=(\tau_{1}\pm i\tau_{2})/2$ and $\Delta_{\rm c(q)}$ is the classical (quantum) component of the SC gap. The check symbol $\check{(\cdot)}$ highlights that the objects are $\times 4$ 4 matrices. The terms in the form $\gamma_{\mu}\check{A}$ stand for the expression $\gamma_{\mu}\otimes\check{A}$, where $\otimes$ is a tensor product. We assume that the SC state is a helical state modulating in the $y$-direction and time-independent. In this case, the Cooper pair momentum corresponds to a uniform gauge field along the $y$-axis $A_{y}$, and we set $A_{x}=0$ from now on.
In addition to Eq. (1), we take into account interaction terms of the Cooper and Coulomb channels. The former can be obtained directly by integrating the spatial fluctuation of the gap function from Eq. (1) [18, 28]. The procedure results in two terms in action. The first term $\bar{S}_{\rm HS}$ has the same form as $S_{\rm HS}$ [Eq. (S.1) in Supplemental Material [35]], but the spatial modulation in the gap functions is removed. Thus, in the following, $\Delta_{\rm c(q)}$ denotes the spatially uniform gap function. The second term is
$S^{\rm(c)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm c}\!\!\sum\limits_{i=1,2}\!\int_{t,\bm{r}}\!\!\!\mathrm{tr}\!\left\{{\gamma_{\rm c}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!\mathrm{tr}\!\left\{{\gamma_{\rm q}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!,$ (2)
[…]
$S^{\rm(s)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm s}\int_{t,\bm{r}}\mathrm{tr}\left\{{\gamma_{\rm c}\tau_{0}Q_{tt}(\bm{r})}\right\}\mathrm{tr}\left\{{\gamma_{\rm q}\tau_{0}Q_{tt}(\bm{r})}\right\},$ (3)
where $\tilde{\Gamma}_{\rm s}=-\pi^{2}\nu\Gamma_{\rm s}/16$ and $\Gamma_{\rm s}=1$ in the short-range limit. Thus, the total action in the partition function $Z$ is given by $S_{\rm tot}=\bar{S}_{\rm HS}+S_{\rm N}+S_{\rm T}+S^{\rm(c)}_{\rm int}+S^{\rm(s)}_{\rm int}$.
Effective action: In the ordinary mean-field theory, the system is described by saddle points of action, which are obtained by searching the matrix field $Q=\bar{Q}$ that optimizes the action under the constraints $\bar{Q}^{2}=1$ and $\mathrm{Tr}\,\{\bar{Q}\}=0$. Going beyond the mean-field approximation can be done by integrating out the fluctuations around the saddle points, resulting in the effective action of the gap function. In this work, we consider the Gaussian fluctuations around a SC saddle point $\bar{Q}_{\rm s}$. Since $\bar{Q}_{\rm s}$ must be on the same manifold as a normal saddle point $\bar{Q}_{\rm N}$, there is a matrix $U$ such that $\bar{Q}_{\rm s}=U\bar{Q}_{\rm N}U^{-1}$. The saddle point configuration may be taken as the fermionic action formulated by Keldysh path integration [14]. That is $\bar{Q}=U_{\rm K}\Lambda U_{\rm K}$, where $\Lambda=\mathrm{diag}\,\{\check{\Lambda}_{\rm R},\check{\Lambda}_{\rm A}\}$ and $\rm R(A)$ denotes the retarded (advanced) component. In the energy representation, $U_{\rm K}$ is
$U_{\mathrm{K}\epsilon}$ $=\begin{pmatrix}\tau_{0}&F_{\epsilon}\tau_{0}\\ 0&-\tau_{0}\end{pmatrix},$ (4)
where distribution in thermal equilibrium leads to $F_{\epsilon}=\tanh\epsilon/(2T)$ with energy $\epsilon$ and temperature $T$. Using the above equations, we can rewrite $\bar{Q}_{\rm N(s)}=U_{\rm K}\Lambda_{\rm N(s)}U_{\rm K}$, where $\Lambda_{\rm N(s)}=\mathrm{diag}\,\{\check{\Lambda}_{\rm N(s),R},\check{\Lambda}_{\rm N(s),A}\}$. The form of $\Lambda_{\rm N}$ in the energy representation is $\Lambda_{\mathrm{N}\epsilon}=\mathrm{diag}\left\{{\tau_{3},-\tau_{3}}\right\}$ [14, 15]. On the other hand, $\Lambda_{\rm s}$ is generally a function of the spectral angle $\theta$, which will be specified later. Note that $U$ is also a function of $\theta$, just as $\Lambda_{\rm s}$ is.
Next, we perform the transformation $Q=U\tilde{Q}U^{-1}$ and then the transformation $\tilde{Q}=\exp\left({W/2}\right)\Lambda_{\rm N}\exp\left({-W/2}\right)$. Thus, $Q$ can be expanded to the quadratic order of $W$ as $Q=\Lambda_{\rm N}+\frac{1}{2}\,[W,\Lambda_{\rm N}]+\frac{1}{8}\,[W,[W,\Lambda_{\rm N}]]$. Expanding the total action to the quadratic order of $W$ results in $S_{\mathrm{tot}}[\Delta,Q]\rightarrow S_{\mathrm{tot}}[\Delta,\theta,W]$, where
$S_{\mathrm{tot}}[\Delta,\theta,W]$ $=S_{0}[W]+S_{\rm MF}[\theta,\Delta]+\tilde{S}_{\rm int}[\theta,W]$
$\quad+S_{\rm I}[\theta,W]+S_{\rm II}[\theta,\Delta,W].$ (5)
Here, $S_{0}[W]$ is the action of the free cooperon and diffuson modes, which are determined from the condition $\{W,\Lambda_{\rm N}\}=0$. The explicit expressions for each term except the last term in Eq. (5) are listed in [35]. The last term $S_{\rm II}$ contains the terms in the form $\sim\mathrm{Tr}\,\{(\dots)WW\}$ and $\sim\mathrm{Tr}\,\{W(\dots)W\}$, which do not make any contribution in the first order of $\Delta_{\rm q}$ and thus are neglected.
Now, we can find the effective action of the gap function, $S_{\rm eff}$. Regarding the terms other than $S_{0}[W]$ as perturbative terms, $S_{\rm eff}$ is given as $\exp\left({iS_{\rm eff}}\right)=\braket{\exp\left({i\left[{S_{\rm MF}+S_{\rm int}+S_{\rm I}}\right]}\right)}_{0}$, where $\braket{\cdots}_{0}$ denotes the average with respect to $S_{0}$. Performing the cumulant expansion, at the one-loop level, the effective action is obtained from the following equation:
$S_{\rm eff}=\braket{S_{\rm MF}+\tilde{S}_{\rm int}}_{0}+\frac{i}{2}\braket{S^{2}_{\rm I}}_{0}.$ (6)
SC saddle point: The SC saddle points $\bar{Q}_{\rm s}$ are usually obtained by numerical techniques. However, since we aim to integrate out the fluctuations to obtain an effective action of the SC gap function, saddle points can be approximated so that all the fluctuations can be considered. Although taking into account all fluctuations is a difficult task, a method can give the familiar GL action for conventional dirty superconductors by calculations at the tree level [16, 17]. Here, we choose the time-independent SC saddle point $\bar{Q}_{\rm s}$ that includes the Zeeman coupling but excludes the interaction terms, the SOC, and the Cooper pair momentum. The saddle point $\bar{Q}_{\rm s}$ is obtained by solving the equation $[\check{E}_{\epsilon}+\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q},\bar{Q}_{\mathrm{s},\epsilon\epsilon}]=0$, where $\check{E}_{\epsilon}=\tau_{3}(\epsilon+h\sigma_{3})$ and $h$ is a Zeeman field. The equation can be easily solved if $\Delta_{\rm q}=0$. However, in the absence of the quantum part $\Delta_{\rm q}$, the solution cannot include any effect of the terms other than the term associated with $\check{E}$ on the effective action. We overcome this problem by instead using the solution of the following equation:
$[\check{E}_{\epsilon}+\check{\Delta}_{\mathrm{R(A)},\epsilon}$ $,\check{\Lambda}_{\mathrm{sR(A)},\epsilon\epsilon}]=0,$ (7)
$\check{\Delta}_{\mathrm{R(A)},\epsilon}=$ $\check{\Delta}_{\rm c}\pm F_{\epsilon}\check{\Delta}_{\rm q}.$ (8)
Solving Eq. (7), we obtain the saddle point $\bar{Q}_{\rm s}=U_{\rm K}\Lambda_{\rm s}U_{\rm K}$ with $\Lambda_{\rm s}=\mathrm{diag}\,\{\check{\Lambda}_{\rm sR},\check{\Lambda}_{\rm sA}\}$, which can be written in the form $\check{\Lambda}_{\rm sR(A)}=\check{U}_{\rm sR(A)}\tau_{3}\check{U}^{-1}_{\rm sR(A)}$. By some calculations, we obtain $\check{U}_{\rm sR(A)}=1-\tau_{1}\hat{X}_{\rm R(A)}$, where
$\hat{X}_{\rm R}=\mathrm{diag}\,$$\Big{\{}\tanh\frac{\theta_{\mathrm{R},\epsilon+}}{2},\tanh\frac{\theta_{\mathrm{R},\epsilon-}}{2}\Big{\}},$ (9)
$\tanh\theta_{\mathrm{R},\epsilon\pm}$ $=\mathrm{sgn}(\epsilon_{\pm})\Delta_{\mathrm{R},\epsilon}/|\epsilon_{\pm}\!+\!i0|,$ (10)
and $\epsilon_{\pm}=\epsilon\pm h$. $\hat{X}_{\rm A}$ has the same form as $\hat{X}_{\rm R}$ but $\theta_{\mathrm{R},\epsilon\pm}$ is replaced by $\theta_{\mathrm{A},\epsilon\pm}$, which satisfies $\tanh\theta_{\mathrm{A},\epsilon\pm}=\Delta_{\mathrm{A},\epsilon}/|\epsilon_{\pm}\!-\!i0|$. This is indeed a possible form of the spectral angle mentioned below Eq. (4). Defining $U\equiv U_{\rm K}U_{\rm s}$ and $U_{\rm s}\equiv\mathrm{diag}\left\{{\check{U}_{\rm sR},\check{U}_{\rm sA}}\right\}$, we can rewritte $\bar{Q}_{\rm s}$ as $\bar{Q}_{\rm s}=U\Lambda_{\rm N}U^{-1}$.
The solution of Eq. (7) can be used to obtain the effective action including the contributions of the SOC, Cooper-pair momentum, and interactions. However, we will go to a more accurate approximation by allowing the spectral angles $\theta_{\mathrm{R(A)},\epsilon\pm}$ to be determined by optimizing the total effective action with respect to them. The outcome is the so-called modified Usadel equation, first proposed in Ref. [28] by the replica formalism. To derive a similar equation in the Keldysh functional formalism, we have to take into account the quantum part of the spectral angle. Otherwise, for the same reason as described in the previous paragraph, the effective action becomes trivial. The problem is how the quantum part of the spectral angles should enter $\Lambda_{\rm s}$. By observing Eqs. (8) and (10), we arrive at the ansatz that the generalized spectral angles should have the form like Eq. (8) for the gap function, and the spectral angles in Eq. (9) should be replaced as follows:
$\theta_{\mathrm{R},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{R},\epsilon\pm}=\theta^{\rm c}_{\mathrm{R},\epsilon\pm}\!\!+F_{\epsilon}\theta^{\rm q}_{\mathrm{R},\epsilon\pm},$ (11)
$\theta_{\mathrm{A},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{A},\epsilon\pm}=\theta^{\rm c}_{\mathrm{A},\epsilon\pm}\!\!-F_{\epsilon}\theta^{\rm q}_{\mathrm{A},\epsilon\pm},$ (12)
where $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$ represent the classical and quantum parts of the generalized spectral angles, respectively. In the above equation, $F_{\epsilon}$ plays an essential role. The effective action would become trivial if there were no term that couples $F_{\epsilon}$ to $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$. Hereafter, we adopt the SC saddle point $\bar{Q}_{\rm s}$ that is determined by Eq. (9) with Eqs. (11) and (12). The quantum parts will be set to zero after performing the first derivatives with respect to them.
Modified Usadel equation: Following the plan described in the previous paragraph, we determine the form of $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ by optimizing the effective action $S_{\rm eff}$ with respect to each spectral angle. We work within GL theory by expanding $\theta_{\mathrm{R(A)},\epsilon\pm}$ as $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}=\Delta_{\rm c}f^{\rm R(A)}_{1\epsilon\pm}+\Delta^{3}_{\rm c}f^{\rm R(A)}_{3\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}=\theta^{\rm R(A)q}_{1\epsilon\pm}+\theta^{\rm R(A)q}_{3\epsilon\pm}$. In this way, $f^{\rm R(A)}_{j\epsilon\pm}$ is determined by the following equation:
$[\delta S_{\rm{eff}}/\delta\theta^{\mu\mathrm{q}}_{j\epsilon\pm}]_{\theta^{\mathrm{q}}\rightarrow 0}=0,$ (13)
where $\mu=\rm R,A$ and $j=1$ or 333. In the following, $f^{\rm R(A)}_{j\epsilon\pm}$ is referred to as spectral function. In the calculation of Eq. (13), the retarded and advanced parts are not independent. Otherwise, irregular divergences appear in the left-hand side of Eq. (13). This argument is explained in detail in Supplemental Material [35]. We find that all irregular terms vanish if the following condition holds:
$f^{\rm R}_{j\epsilon\pm}=-f^{\rm A}_{j-\epsilon\mp},\quad\theta^{\rm Rq}_{j\epsilon\pm}=-\theta^{\rm Aq}_{j-\epsilon\mp}.$ (14)
Using Eqs. (11)--(14) and performing analytic continuation, we obtain the modified Usadel equations at the first and third orders of $\Delta_{\rm c}$ as follows:
$\mathcal{S}_{\mathrm{kernel},j}[f_{1n\pm},f_{3n\pm}]=\mathcal{S}_{\mathrm{source},j}[f_{1n\pm},f_{3n\pm}],$ (15)
where $f_{jn\pm}=if^{\rm R}_{ji\omega_{n}\pm}$, $\omega_{n\pm}=\omega_{n}\pm ih$, $\omega_{n}=(2n+1)\pi T~(n\in\mathbb{Z})$, and $\mathcal{S}_{\mathrm{source},1}=1$. The others are shown in Supplemental Material [35] to avoid describing the lengthy equations, and we find that the Cooper interaction terms are not involved in the equation determining $f_{1n\pm}$. The equations are solved numerically by the Padé decomposition method [36], which can achieve fast convergence when calculating the Matsubara summation.
Equation of state: The gap equation is obtained from the stationary condition $[\partial S/\partial\Delta_{\rm q}]_{\Delta_{\rm q}\rightarrow 0}=0$. In GL theory, the equation has the form $\alpha_{2}\Delta_{\rm c}+\alpha_{4}\Delta_{\rm c}^{3}=0$, where
$\alpha_{2}$ $=4\,\Gamma^{-1}_{\rm c}\!-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{[}f_{1n\lambda}\!+\!\frac{1}{\omega_{n}\!+\!i\lambda h}\Big{]}$
$\quad-2\int_{0}^{\omega_{\rm D}}\frac{d\epsilon}{\epsilon}\Big{[}\tanh\left({\frac{\epsilon_{+}}{2T}}\right)+\tanh\left({\frac{\epsilon_{-}}{2T}}\right)\Big{]},$ (16)
$\alpha_{4}$ $=-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{(}f_{3n\lambda}-\!\frac{1}{6}[f_{1n\lambda}]^{3}\Big{)}.$ (17)
Note that $f_{jn\lambda}$ is derived from Eq. (15), and $\alpha_{2}$ is regularized by introducing the ultraviolet cutoff, which is the Debye energy $\omega_{\rm D}$, to the energy integration. The form of the gap equation in the GL level suggests that the free energy and the supercurrent are given as $f_{\mathrm{GL}}=\frac{1}{2}\alpha_{2}\Delta^{2}_{\rm c}+\frac{1}{4}\alpha_{4}\Delta^{4}_{\rm c}$ and $I_{\rm s}=\partial f_{\mathrm{GL}}/\partial A_{y}$, respectively. The helical SC state for any supercurrent $I_{s}$ is then determined by the latter equation. In particular, the second-order SC phase transition line in the absence of the supercurrent is calculated based on the conditions $\alpha_{2}=0$ and $\partial\alpha_{2}/\partial A_{y}=0$. From now on, let $A_{0}$ be the $A_{y}$ field in the zero-current SC states. We denote the transition temperature at a magnetic field $h$ as $T_{\rm c}(h)$, which is denoted as $T_{\rm c0}(h)$ when we neglect the e-e interaction. At $h=0$, we describe $T_{\rm c0}(0)=T_{\rm c0}$ and $T_{\rm c}(0)=T_{\rm c}$, for simplicity. For later use, we define the reduced temperature $\tau_{\rm r}$ as $\tau_{\rm r}=T/T_{\rm c0}$ or $\tau_{\rm r}=T/T_{\rm c}$.
WL conductivity: When the current reaches its critical value, the SC state changes to a resistive state. Here, we study the WL correction to the conductivity, $\sigma_{\rm WL}$ [37, 38, 39]. For simplicity, we neglect the e-e interactions. We find that it is convenient to work with the alternative normal saddle point $\bar{Q}^{\prime}_{\rm N}=U^{\prime}\Lambda^{\prime}_{\rm N}U^{\prime-1}$, where $\Lambda^{\prime}_{\rm N}=\mathrm{diag}\,\{\gamma_{3},\gamma_{3}\}$ and $U^{\prime}=\mathrm{diag}\,\{\check{U}_{\rm K},\check{U}^{\mathbb{T}}_{\rm K}\}$. Here, $U^{\mathbb{T}}$ denotes the transpose of a matrix and the interchange of the time of $U$, which means that $U^{\mathbb{T}}_{t_{1}t_{2}}=U^{\rm t}_{t_{2}t_{1}}$ with $U^{\rm t}_{t_{2}t_{1}}$ being the ordinary transpose in matrix elements of $U_{t_{2}t_{1}}$. The matrix $\check{U}_{\rm K}$ has the same form as Eq. (4), but $\tau_{0}$ is replaced by $\sigma_{0}$.
This saddle point comes from the time-reversal symmetry condition of the matrix field $Q$. To the best of our knowledge, this saddle point was first introduced in Ref. 40 to calculate the WL conductivity $\sigma_{\rm WL}$ in the spinless case within the Keldysh functional formalism. Here, we generalize their formalism to include spin degrees of freedom. We explain the procedure in [35], and the equation for $\sigma_{\rm WL}$ is consistent with existing results [41, 42, 44, 43].
Numerical results: In the numerical calculations, we fix the parameters, $\Gamma_{\rm c}=0.2$, $\omega_{\rm D}/T_{\rm c0}=0.1309$, and Fermi energy $E_{\rm F}/T_{\rm c0}=p_{\rm F}^{2}/(2mT_{\rm c0})=10^{3}$, and consider the following two sets of parameters: (I) $\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$ and (II) $\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$, where $\alpha$ is the Rashba SOC constant and $\tau$ is the impurity scattering time. Note that the values of $\tau T_{\rm c0}$ in cases (I) and (II) correspond to the characteristic lengths of $l_{\rm F}p_{\rm F}=30$ and 151515, respectively, which are consistent with the WL regime. The characteristic energies of SOC and disorder are represented by $\epsilon_{\rm SOC}=\alpha v_{\rm F}$ and $\epsilon_{\rm D}=1/\tau$, respectively. In case (I), the SOC is comparable with disorder as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.6$, while case (II) is in a weak SOC regime as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.06$. We show the numerical results for the comparable and weak SOC in Figs. 1 and 2, respectively.
Figure 1: The diode quality factor $\eta$ and the WL correction to conductivity $\sigma_{\rm WL}$ in the intermediate SOC regime ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$). The solid (dashed) lines show the results in the interacting (free) case with (without) the e-e interaction. Panels (a) and (c) show the color maps of $\sigma_{\rm WL}$ and $\eta$, respectively. The solid and dashed lines show the second-order superconducting transition line. The bold circles in (a) represent the tricritical points. Panels (b) and (d) show $\sigma_{\rm WL}$ and $\eta$ near the SC transition line, respectively. In these figures, we vary the reduced temperature $\tau_{\rm r}$ and magnetic field $h$ so that the temperature is $T=0.99T_{\rm c}(h)$.
We first show the properties of zero-current SC states. As shown in Figs. 1a and 2a, the e-e interactions suppress the transition temperatures and critical magnetic fields. The gauge field $A_{0}$ corresponding to the momentum of Cooper pairs is also suppressed when the interactions are incorporated [35]. Since there is no contribution from the Cooper channel in the equation determining the second-order SC transition line, the suppression of the transition temperature is attributed to the Coulomb channel. The zero-field transition temperatures are $T_{\rm c}=0.8793T_{\rm c0}$ and $T_{\rm c}=0.6457T_{\rm c0}$ in Figs. 1 and 2, respectively. We have confirmed that $T_{\rm c}$ decreases as the impurity scattering time $\tau$ decreases, while it increases toward $T_{\rm c0}$ as the SOC $\alpha$ increases. The transition lines depicted in Figs. 1a and 2a show the second-order phase transition, and they terminate at the tricritical points at high magnetic fields, indicating the first-order transition, at which $\alpha_{4}$ becomes negative. It is well known that the first-order SC transition occurs in the high field region due to the Pauli paramagnetic effects when we adopt the mean-field theory for spatially uniform $s$-wave states [45]. At magnetic fields higher than the tricritical point, the second-order transition from the normal state to the non-uniform stripe SC state is expected [46]. Since the Cooper channel makes a non-zero contribution to $\alpha_{4}$, both Coulomb and Cooper channels affect the tricritical points.
Figure 2: The WL correction to conductivity $\sigma_{\rm WL}$ in the weak SOC regime ($\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$). Panel (a) shows the color map of $\sigma_{\rm WL}$, which is plotted in the same manner as Fig. 1a. Panel (b) shows $\sigma_{\rm WL}$ near the superconducting transition line in the same manner as Fig. 1b.
Next, let us focus on the quality factor $\eta$ of the SD effect. The diode quality factor $\eta$ is defined as $\eta=(1-\xi)/(1+\xi)$, where $\xi=|I_{\rm sc-}/I_{\rm sc+}|$ and $I_{\rm sc\pm}$ is the critical current in the ±$\pm y$-direction. Because $\eta$ is tiny in the weak SOC regime, we calculate the SDE in the intermediate SOC regime and show the results of $\eta$ in Figs. 1c and 1d. As shown in Fig. 1c, the diode quality factor increases as the system approaches the tricritical point. To show the behaviors of the SDE near the transition line, we plot in Fig. 1d the diode quality factor near the transition line as a function of the reduced temperature. The absolute values of the diode quality factor increase divergently as the tricritical point is approached. We see qualitatively the same behaviors in the absence and presence of the e-e interaction, while the tricritical point is shifted. Although the Coulomb interaction significantly reduces the zero-field transition temperature, the diode quality factor almost coincides in both cases at high reduced temperatures. This behavior is deviated as the tricritical points are approached. Because the tricritical point is shifted by interactions, the absolute values of the diode quality factor in the presence of interactions become smaller than those in the case without interactions.
Finally, we discuss the WL correction to the conductivity $\sigma_{\rm WL}$. The correction is normalized with the Drude conductivity $\sigma_{\rm D}=2\nu D$. As shown in Fig. 1a, interestingly, $\sigma_{\rm WL}$ in the intermediate SOC regime decreases as the temperature and magnetic field increase, indicating stronger localization effects. By examining the plot near the transition lines [Fig. 1b], we conclude that the e-e interaction can suppress the WL correction. However, as shown in Fig. 2a, the qualitatively opposite behaviors manifest themselves in almost the entire region of temperatures and magnetic fields. Looking at the results near the transition lines [Fig. 2b], we find that the e-e interaction can enhance the WL correction when the SOC is weak.
Summary: We have formulated the renormalized GL theory including the Cooper and Coulomb interaction terms from the NLSM based on the Keldysh functional formalism. Our formalism extends the previous one and enables us to study dirty 2D Rashba superconductors under in-plane Zeeman fields, which are typical platforms of the SD effect. We find that the interplay between localization and interactions leads to the corrections that suppress the SC states, the first-order transition, and the SD effect. Furthermore, we have calculated the WL correction to conductivity that characterizes transport when the SC state is destroyed by an electric current exceeding the critical current. The WL correction can be suppressed or enhanced by interactions depending on the strength of SOC. We expect our results to pave the way for switching between superconducting, metallic, and insulating transport via an electric current.
Acknowledgements — We appreciate helpful discussions with Ryusuke Ikeda, Hideki Narita, Teruo Ono, and Ryuji Hakuno. This work was supported by JSPS KAKENHI (Grant Numbers JP22H01181, JP22H04933, JP23K17353, JP23K22452, JP24K21530, JP24H00007, JP25H01249).

### v2

@@
 Abstract
-We study a dirty two-dimensional superconductor with Rashba spin-orbit coupling and in-plane Zeeman fields described by the nonlinear sigma model that includes short-range electron-electron interactions from the Coulomb and Cooper channels.
-The renormalized Ginzburg-Landau theory, which includes the weak localization effects at the one-loop level, is constructed by using the Keldysh functional formalism.
-It is shown that the tricritical point appears in the phase diagram.
-The superconducting diode quality factor increases divergently as the system approaches the tricritical point.
-Near the superconducting phase transition lines, the absolute value of the diode quality factor decreases due to the cooperation of localization and interactions.
-The normal conductivity of the resistive state, in which the superconducting state is destroyed by the critical current, is calculated, and localization behaviors are demonstrated.
+We study a dirty two-dimensional superconductor with Rashba spin-orbit coupling and in-plane Zeeman fields described by the nonlinear sigma model that includes the Cooper and long-range Coulomb interactions.
+The renormalized Ginzburg-Landau theory, which includes the weak localization effects at the one-loop level, is constructed using the Keldysh functional formalism.
+It is shown that the transition temperature and magnetic field, as well as the tricritical point appearing in the phase diagram, are suppressed by the interactions.
+Nevertheless, we have found a universal behavior in the high-transition-temperature regime that demonstrates the robustness of the superconducting diode effect against the interactions.
+The conductivity of the resistive states emerging after the superconducting states are destroyed by the critical current is also calculated, and localization behaviors are demonstrated.
 Introduction: The phenomenon in which the critical currents of superconductors become nonreciprocal, known as the superconducting diode (SD) effect [1, 2], has attracted much attention.
 Theoretical studies have been carried out for various types of noncentrosymmetric superconductors [3, 4, 5, 6, 7, 8, 9].
-Not only clean systems but also systems with uncorrelated impurities have been studied in detail at the mean-field level [10].
+Not only clean systems but also disordered systems have been studied at the mean-field (MF) level [10, 11, 12].
 If electron systems are disordered enough, they are well described by the nonlinear sigma model (NLSM) [13, 14, 15, 16], and dirty superconductors can also be described by the NLSM [17, 18, 19, 20].
-Recently, an NLSM for a dirty superconductor with spin-orbit coupling (SOC) and a Zeeman field was derived [19] and used to investigate the SD effect in Josephson junctions at the mean-field level [20].
-However, the Anderson localization effect, a correction to the mean-field theory in dirty systems, has not yet been highlighted in studies of the SD effect.
+Recently, an NLSM for a dirty superconductor with spin-orbit coupling (SOC) and Zeeman field was derived [21] and used to investigate the SD effect in Josephson junctions at the MF level [22].
+However, the Anderson localization effect, a correction to the MF theory in dirty systems, has not yet been highlighted in studies of the SD effect.
 The role of Anderson localization in superconductors is an important topic of ongoing debate.
-Without considering strong mesoscopic correlations and electron-electron (e-e) interactions, the properties of dirty $s$-wave superconductors at zero magnetic field remain unchanged with respect to the strength of non-magnetic disorder in the weak localization (WL) regime, known as the Anderson theorem [21].
-If the Coulomb interaction is concerned, the transition temperature for the $s$-wave pairing states is suppressed in the WL regime [22, 23, 24].
-The opposite effect can occur in the strong localization regime by the multifractality effect [25, 26, 27, 28].
-It was also shown that, in the WL regime, the Coulomb interaction suppresses the critical current of conventional dirty superconductors at zero magnetic field and low temperatures [29, 30].
-In this Letter, we study the interplay of the e-e interaction and disorder on the SD effect and clarify the WL correction to conductivity when the superconducting (SC) state is broken by the critical current.
-We work with the NLSM for a dirty two-dimensional (2D) noncentrosymmetric superconductor, including the interactions from the Coulomb and Cooper channels in the short-range limit, and establish a renormalized Ginzburg-Landau (GL) theory at the one-loop level.
+Without electron-electron (e-e) interactions, the properties of dirty $s$-wave superconductors at zero magnetic field remain unchanged with respect to the strength of non-magnetic disorder in the weak localization (WL) regime, known as the Anderson theorem [23].
+However, if the long-range Coulomb interaction exists, the transition temperature is largely suppressed even in the WL regime [24, 25, 26].
+Moreover, the transition temperature can also be enhanced by the multifractality effect [27, 28, 29, 30], where the short-range nature of the e-e interactions is important.
+It was also shown that, in the WL regime, the Coulomb interaction lowers the critical current of conventional dirty superconductors at zero magnetic field and low temperatures [31, 32].
+In this Letter, we study the role of WL correction from the Cooper and long-range Coulomb interactions, which are ubiquitous in the solid state systems [33, 34, 35, 36, 37], on the SD effect.
+In normal metals, the localization behavior can be changed by the SOC.
+It is interesting to consider whether the WL behavior of the resistive states that emerge after the superconducting (SC) states are destroyed by the critical currents can be enhanced while maintaining the high efficiency of the SD effect.
+We thus also investigate the relationship between the localization behavior in such resistive states and the SD effect.
+We work with the NLSM for a dirty two-dimensional (2D) noncentrosymmetric superconductor, and establish a renormalized Ginzburg-Landau (GL) theory at the one-loop level.
 While many studies [30, 38, 39, 40] have been conducted in the replica formalism, we formulate the theory with the Keldysh functional formalism [16, 17, 19, 41].
-Model: We consider a dirty 2D superconductor aligned in the $xy$-plane with the Rashba SOC and a Zeeman field in the $x$-direction.
-In the Keldysh functional formalism, the partition function is given as $Z=\int D[Q,\Delta]\exp\left({iS[Q,\Delta]}\right)$ with [19]
-$S[\Delta,Q]$ $=S_{\rm HS}[\Delta]+S_{\rm N}[\Delta,Q]+S_{\rm T}[Q],$ (1)
-where $S_{\rm HS}$ is responsible for the Hubbard-Stratonovich transformation associated with the SC state, $S_{\rm N}$ is the conventional non-interacting part, and $S_{\rm T}$ is the topological part.
-Their explicit forms are shown in Supplemental material [35].
-A $\times 8$ 8 matrix field $Q$ satisfies the conditions $Q^{2}=1$ and $\mathrm{Tr}\,\{Q\}=0$.
-Let $\gamma_{\mu},~\tau_{\mu}$, and $\sigma_{\mu}~(\mu=0,\dots,4)$ be the Pauli matrices in the Keldysh, Nambu, and spin spaces, respectively.
-In particular, we write $\gamma_{\rm c}=\gamma_{0}$ and $\gamma_{\rm q}=\gamma_{1}$ and often abbreviate $\gamma_{0}$ for simplicity.
-The order parameter of superconductivity, that is, the SC gap, is then described by the matrix $\gamma_{\rm c}\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q}$, where $\check{\Delta}_{\rm c(q)}=\tau_{+}\Delta_{\rm c(q)}+\tau_{-}\Delta^{*}_{\rm c(q)}$ with $\tau_{\pm}=(\tau_{1}\pm i\tau_{2})/2$ and $\Delta_{\rm c(q)}$ is the classical (quantum) component of the SC gap.
-The check symbol $\check{(\cdot)}$ highlights that the objects are $\times 4$ 4 matrices.
-The terms in the form $\gamma_{\mu}\check{A}$ stand for the expression $\gamma_{\mu}\otimes\check{A}$, where $\otimes$ is a tensor product.
-We assume that the SC state is a helical state modulating in the $y$-direction and time-independent.
-In this case, the Cooper pair momentum corresponds to a uniform gauge field along the $y$-axis $A_{y}$, and we set $A_{x}=0$ from now on.
+Model: We consider a dirty 2D superconductor aligned in the $xy$-plane with the Rashba SOC and a Zeeman field $h$ in the $x$-direction.
+It was shown in [21] that in the Keldysh functional formalism, the partition function of the system without e-e interactions is given as $Z=\int D[Q,\Delta]\exp\left({iS[Q,\Delta]}\right)$, where
+$S[\Delta,Q]$ $=S_{\rm HS}[\Delta]+S_{\rm N}[\Delta,Q]+S_{\rm T}[Q].$ (1)
+[…]
+In Eq.
+(1), the fluctuation of the gap matrix $\Delta$ from the static helical SC state with a fixed Cooper pair momentum $\bm{P}$ is already integrated out.
+Here, the static helical SC state is characterized by the gap function $\Delta_{\rm cl(q)}\exp\,(i\bm{P}\cdot\bm{r})$, where $\Delta_{\rm cl(q)}\in\mathbb{R}$ is the classical (quantum) component of the SC gap, and it does not depend on space and time.
+See the procedure in [42].
+Thus, only the field $Q$ and the static uniform SC gaps remain in the integral measure of the partition function $Z$.
+Moreover, since we consider the Zeeman field $h$ pointing in the $x$-direction, the SD effect occurs when a supercurrent flows in the $y$-direction.
+We thus set $P_{x}=0$.
+The factor $\exp\,(iP_{y}y)$ can also be cancelled out by a U(1) transformation.
+As a result, in Eq.
+(1), a uniform gauge field, $A_{y}=P_{y}/2$, appears in the covariant derivative, and $\Delta$ is expressed as $\Delta=\gamma_{\rm cl}\check{\Delta}_{\rm cl}+\gamma_{\rm q}\check{\Delta}_{\rm q}$, where $\check{\Delta}_{\rm cl(q)}=i\tau_{2}\Delta_{\rm cl(q)}$.
 In addition to Eq.
 (1), we take into account interaction terms of the Cooper and Coulomb channels.
-The former can be obtained directly by integrating the spatial fluctuation of the gap function from Eq.
-(1) [18, 28].
-The procedure results in two terms in action.
-The first term $\bar{S}_{\rm HS}$ has the same form as $S_{\rm HS}$ [Eq.
-(S.1) in Supplemental Material [35]], but the spatial modulation in the gap functions is removed.
-Thus, in the following, $\Delta_{\rm c(q)}$ denotes the spatially uniform gap function.
-The second term is
-$S^{\rm(c)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm c}\!\!\sum\limits_{i=1,2}\!\int_{t,\bm{r}}\!\!\!\mathrm{tr}\!\left\{{\gamma_{\rm c}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!\mathrm{tr}\!\left\{{\gamma_{\rm q}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!,$ (2)
-[…]
-$S^{\rm(s)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm s}\int_{t,\bm{r}}\mathrm{tr}\left\{{\gamma_{\rm c}\tau_{0}Q_{tt}(\bm{r})}\right\}\mathrm{tr}\left\{{\gamma_{\rm q}\tau_{0}Q_{tt}(\bm{r})}\right\},$ (3)
-where $\tilde{\Gamma}_{\rm s}=-\pi^{2}\nu\Gamma_{\rm s}/16$ and $\Gamma_{\rm s}=1$ in the short-range limit.
-Thus, the total action in the partition function $Z$ is given by $S_{\rm tot}=\bar{S}_{\rm HS}+S_{\rm N}+S_{\rm T}+S^{\rm(c)}_{\rm int}+S^{\rm(s)}_{\rm int}$.
-Effective action: In the ordinary mean-field theory, the system is described by saddle points of action, which are obtained by searching the matrix field $Q=\bar{Q}$ that optimizes the action under the constraints $\bar{Q}^{2}=1$ and $\mathrm{Tr}\,\{\bar{Q}\}=0$.
-Going beyond the mean-field approximation can be done by integrating out the fluctuations around the saddle points, resulting in the effective action of the gap function.
-In this work, we consider the Gaussian fluctuations around a SC saddle point $\bar{Q}_{\rm s}$.
-Since $\bar{Q}_{\rm s}$ must be on the same manifold as a normal saddle point $\bar{Q}_{\rm N}$, there is a matrix $U$ such that $\bar{Q}_{\rm s}=U\bar{Q}_{\rm N}U^{-1}$.
-The saddle point configuration may be taken as the fermionic action formulated by Keldysh path integration [14].
+Let $S^{\rm(c)}_{\rm int}$ and $S^{\rm(s)}_{\rm int}$ denote the Cooper and Coulomb interaction actions, respectively.
+Their explicit forms are shown in [42].
+Note that we consider the Coulomb interaction in the long-range limit.
+The total partition function $Z_{\rm tot}$ is given as $Z_{\rm tot}=\int D[Q,\Delta]\exp\left({iS_{\rm tot}[Q,\Delta]}\right)$, where
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}S_{\rm tot}[\Delta,Q]=S[\Delta,Q]+S^{\rm(c)}_{\rm int}[Q]+S^{\rm(s)}_{\rm int}[Q].$} (2)
+Despite being beyond our scope, it is instructive to note that the triplet channel of e-e interactions can be neglected in any case if the triplet diffusive modes acquire a large mass.
+The diffusive modes, which are cooperon and diffuson modes, will be described later.
+In our case, the mass of the triplet diffusive modes is the DP relaxation rate $\Gamma_{\rm r}$.
+Further details are discussed in [42].
+Effective action: In the NLSM, the system is described at the MF level by searching the saddle matrix field $Q=\bar{Q}$ that optimizes the action under the constraints $\bar{Q}^{2}=1$ and $\mathrm{Tr}\,\{\bar{Q}\}=0$.
+It is possible to optimize the total action $S_{\rm tot}$ to obtain an MF theory that includes the contribution of e-e interactions [30].
+However, the corresponding Usadel equation has a nonlinear form that is difficult to solve.
+We thus take an alternative track [30]: Setting a saddle point and integrating out fluctuations around it.
+This ultimately leads to an effective action of the SC gap.
+The procedure to achieve such effective action is outlined below.
+We consider the Gaussian fluctuations around an SC saddle point $\bar{Q}_{\rm s}$ whose form will be specified later.
+Although it is possible to adopt the normal saddle point $\bar{Q}_{\rm N}$ as the origin of expansion, the high-order calculation of perturbation theory is needed [19].
+Time and effort can be saved by considering the SC saddle point [18].
+Since $\bar{Q}_{\rm N}$ and $\bar{Q}_{\rm s}$ satisfy $\bar{Q}^{2}_{\rm N}=\bar{Q}^{2}_{\rm s}=1$, there is a matrix $U$ such that $\bar{Q}_{\rm s}=U\bar{Q}_{\rm N}U^{-1}$.
+The saddle point configuration is the same as the Keldysh fermionic action [16].
 That is $\bar{Q}=U_{\rm K}\Lambda U_{\rm K}$, where $\Lambda=\mathrm{diag}\,\{\check{\Lambda}_{\rm R},\check{\Lambda}_{\rm A}\}$ and $\rm R(A)$ denotes the retarded (advanced) component.
@@
 Next, we perform the transformation $Q=U\tilde{Q}U^{-1}$ and then the transformation $\tilde{Q}=\exp\left({W/2}\right)\Lambda_{\rm N}\exp\left({-W/2}\right)$.
-Thus, $Q$ can be expanded to the quadratic order of $W$ as $Q=\Lambda_{\rm N}+\frac{1}{2}\,[W,\Lambda_{\rm N}]+\frac{1}{8}\,[W,[W,\Lambda_{\rm N}]]$.
+Thus, $Q$ can be expanded to the quadratic order of $W$ as $\tilde{Q}=\Lambda_{\rm N}+\frac{1}{2}\,[W,\Lambda_{\rm N}]+\frac{1}{8}\,[W,[W,\Lambda_{\rm N}]]$.
 Expanding the total action to the quadratic order of $W$ results in $S_{\mathrm{tot}}[\Delta,Q]\rightarrow S_{\mathrm{tot}}[\Delta,\theta,W]$, where
@@
 (4) are listed in [42].
-The last term $S_{\rm II}$ contains the terms in the form $\sim\mathrm{Tr}\,\{(\dots)WW\}$ and $\sim\mathrm{Tr}\,\{W(\dots)W\}$, which do not make any contribution in the first order of $\Delta_{\rm q}$ and thus are neglected.
+The last term $S_{\rm II}$ contains terms in the form $\sim\mathrm{Tr}\,\{(\dots)WW\}$ and $\sim\mathrm{Tr}\,\{W(\dots)W\}$, which do not make any contribution in the first order of $\Delta_{\rm q}$ and thus are neglected.
 Now, we can find the effective action of the gap function, $S_{\rm eff}$.
 Regarding the terms other than $S_{0}[W]$ as perturbative terms, $S_{\rm eff}$ is given as $\exp\left({iS_{\rm eff}}\right)=\braket{\exp\left({i\left[{S_{\rm MF}+S_{\rm int}+S_{\rm I}}\right]}\right)}_{0}$, where $\braket{\cdots}_{0}$ denotes the average with respect to $S_{0}$.
-Performing the cumulant expansion, at the one-loop level, the effective action is obtained from the following equation:
+Performing the cumulant expansion at the one-loop level, the effective action is obtained from the following equation:
 $S_{\rm eff}=\braket{S_{\rm MF}+\tilde{S}_{\rm int}}_{0}+\frac{i}{2}\braket{S^{2}_{\rm I}}_{0}.$ (5)
-SC saddle point: The SC saddle points $\bar{Q}_{\rm s}$ are usually obtained by numerical techniques.
-However, since we aim to integrate out the fluctuations to obtain an effective action of the SC gap function, saddle points can be approximated so that all the fluctuations can be considered.
-Although taking into account all fluctuations is a difficult task, a method can give the familiar GL action for conventional dirty superconductors by calculations at the tree level [16, 17].
-Here, we choose the time-independent SC saddle point $\bar{Q}_{\rm s}$ that includes the Zeeman coupling but excludes the interaction terms, the SOC, and the Cooper pair momentum.
-The saddle point $\bar{Q}_{\rm s}$ is obtained by solving the equation $[\check{E}_{\epsilon}+\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q},\bar{Q}_{\mathrm{s},\epsilon\epsilon}]=0$, where $\check{E}_{\epsilon}=\tau_{3}(\epsilon+h\sigma_{3})$ and $h$ is a Zeeman field.
-The equation can be easily solved if $\Delta_{\rm q}=0$.
-However, in the absence of the quantum part $\Delta_{\rm q}$, the solution cannot include any effect of the terms other than the term associated with $\check{E}$ on the effective action.
-We overcome this problem by instead using the solution of the following equation:
-$[\check{E}_{\epsilon}+\check{\Delta}_{\mathrm{R(A)},\epsilon}$ $,\check{\Lambda}_{\mathrm{sR(A)},\epsilon\epsilon}]=0,$ (7)
-$\check{\Delta}_{\mathrm{R(A)},\epsilon}=$ $\check{\Delta}_{\rm c}\pm F_{\epsilon}\check{\Delta}_{\rm q}.$ (8)
+SC saddle point: Choosing the appropriate saddle point is a key to a good approximation of perturbation theory.
+It is naively expected that the best SC saddle point $\bar{Q}_{\rm s}$ should optimize the effective action Eq.
+(5).
+However, an estimation of $\bar{Q}_{\rm s}$ is necessary to make the analysis possible.
+This is achieved by setting a saddle point and then incorporating contributions from terms unrelated to the saddle point.
+We choose a saddle point that includes the Zeeman coupling.
+In the ordinary MF theory, this is obtained by solving equation $\check{E}_{\epsilon}+\check{\Delta}_{\rm cl}+\gamma_{\rm q}\check{\Delta}_{\rm q},\bar{Q}_{\mathrm{s},\epsilon\epsilon}]=0$ with $\Delta_{\rm q}=0$, where $\check{E}_{\epsilon}=\tau_{3}(\epsilon+h\sigma_{3})$.
+However, the solution does not include any effect of the terms unrelated to the saddle point because the quantum part necessary for deriving the gap equation is absent.
+To overcome this, we demand the retarded and advanced components of the equation to be zero, i.e.,
+$\quad\quad~[\check{E}_{\epsilon}+\check{\Delta}_{\mathrm{R(A)},\epsilon},\check{\Lambda}_{\mathrm{sR(A)},\epsilon\epsilon}]=0,$ (6)
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\check{\Delta}_{\mathrm{R(A)},\epsilon}=i\tau_{2}\Delta_{\mathrm{R(A)},\epsilon}\equiv i\tau_{2}\,(\Delta_{\rm cl}\pm F_{\epsilon}\Delta_{\rm q}).}$ (7)
 Solving Eq.
@@
 (3).
-Defining $U\equiv U_{\rm K}U_{\rm s}$ and $U_{\rm s}\equiv\mathrm{diag}\left\{{\check{U}_{\rm sR},\check{U}_{\rm sA}}\right\}$, we can rewritte $\bar{Q}_{\rm s}$ as $\bar{Q}_{\rm s}=U\Lambda_{\rm N}U^{-1}$.
-The solution of Eq.
-(7) can be used to obtain the effective action including the contributions of the SOC, Cooper-pair momentum, and interactions.
-However, we will go to a more accurate approximation by allowing the spectral angles $\theta_{\mathrm{R(A)},\epsilon\pm}$ to be determined by optimizing the total effective action with respect to them.
-The outcome is the so-called modified Usadel equation, first proposed in Ref.
-[28] by the replica formalism.
-To derive a similar equation in the Keldysh functional formalism, we have to take into account the quantum part of the spectral angle.
-Otherwise, for the same reason as described in the previous paragraph, the effective action becomes trivial.
-The problem is how the quantum part of the spectral angles should enter $\Lambda_{\rm s}$.
-By observing Eqs.
-(8) and (10), we arrive at the ansatz that the generalized spectral angles should have the form like Eq.
-(8) for the gap function, and the spectral angles in Eq.
-(9) should be replaced as follows:
-$\theta_{\mathrm{R},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{R},\epsilon\pm}=\theta^{\rm c}_{\mathrm{R},\epsilon\pm}\!\!+F_{\epsilon}\theta^{\rm q}_{\mathrm{R},\epsilon\pm},$ (11)
-$\theta_{\mathrm{A},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{A},\epsilon\pm}=\theta^{\rm c}_{\mathrm{A},\epsilon\pm}\!\!-F_{\epsilon}\theta^{\rm q}_{\mathrm{A},\epsilon\pm},$ (12)
-where $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$ represent the classical and quantum parts of the generalized spectral angles, respectively.
-In the above equation, $F_{\epsilon}$ plays an essential role.
-The effective action would become trivial if there were no term that couples $F_{\epsilon}$ to $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$.
+Defining $U\equiv U_{\rm K}U_{\rm s}$ and $U_{\rm s}\equiv\mathrm{diag}\left\{{\check{U}_{\rm sR},\check{U}_{\rm sA}}\right\}$, we can rewrite $\bar{Q}_{\rm s}$ as $\bar{Q}_{\rm s}=U\Lambda_{\rm N}U^{-1}$.
+The solution presented above can be used to obtain the effective action incorporating the contributions of the SOC, Cooper-pair momentum, and interactions.
+This is allowed by the presence of $F_{\epsilon}$ in Eq.
+(7).
+Without this term, the effective action would become trivial through the energy integration.
+However, since we search the saddle point that optimizes Eq.
+(5), some terms in Eq.
+(8) need to be specified by the saddle point condition for Eq.
+(5).
+Similarly to Ref.
+[30], we leave the spectral angle as an unknown quality.
+By paying attention to the role of $F_{\epsilon}$ described above, we arrive at the ansatz that the spectral angles in Eq.
+(8) should be replaced by the generalized spectral angles shown below:
+$\theta_{\mathrm{R},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{R},\epsilon\pm}=\theta^{\rm cl}_{\mathrm{R},\epsilon\pm}\!\!+F_{\epsilon}\theta^{\rm q}_{\mathrm{R},\epsilon\pm},$ (10)
+$\theta_{\mathrm{A},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{A},\epsilon\pm}=\theta^{\rm cl}_{\mathrm{A},\epsilon\pm}\!\!-F_{\epsilon}\theta^{\rm q}_{\mathrm{A},\epsilon\pm},$ (11)
+where $\theta^{\rm cl}_{\mathrm{R(A)},\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$ represent the classical and quantum parts of the generalized spectral angles, respectively.
 Hereafter, we adopt the SC saddle point $\bar{Q}_{\rm s}$ that is determined by Eq.
@@
 Modified Usadel equation: Following the plan described in the previous paragraph, we determine the form of $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ by optimizing the effective action $S_{\rm eff}$ with respect to each spectral angle.
-We work within GL theory by expanding $\theta_{\mathrm{R(A)},\epsilon\pm}$ as $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}=\Delta_{\rm c}f^{\rm R(A)}_{1\epsilon\pm}+\Delta^{3}_{\rm c}f^{\rm R(A)}_{3\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}=\theta^{\rm R(A)q}_{1\epsilon\pm}+\theta^{\rm R(A)q}_{3\epsilon\pm}$.
+We work within the GL theory by expanding $\theta_{\mathrm{R(A)},\epsilon\pm}$ as $\theta^{\rm cl}_{\mathrm{R(A)},\epsilon\pm}=\Delta_{\rm cl}f^{\rm R(A)}_{1\epsilon\pm}+\Delta^{3}_{\rm cl}f^{\rm R(A)}_{3\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}=\theta^{\rm R(A)q}_{1\epsilon\pm}+\theta^{\rm R(A)q}_{3\epsilon\pm}$.
 In this way, $f^{\rm R(A)}_{j\epsilon\pm}$ is determined by the following equation:
-$[\delta S_{\rm{eff}}/\delta\theta^{\mu\mathrm{q}}_{j\epsilon\pm}]_{\theta^{\mathrm{q}}\rightarrow 0}=0,$ (13)
-where $\mu=\rm R,A$ and $j=1$ or 333.
+$\delta S_{\rm{eff}}/\delta\theta^{\mu\mathrm{q}}_{j\epsilon\pm}]_{\theta^{\mathrm{q}}\rightarrow 0}=0,$ (12)
+where $\mu=\rm R,A$ and $j=1$ or $3$.
 In the following, $f^{\rm R(A)}_{j\epsilon\pm}$ is referred to as spectral function.
-In the calculation of Eq.
-(13), the retarded and advanced parts are not independent.
+Note that this spectral function $f^{\rm R(A)}_{j\epsilon\pm}$ is different from the Fermi distribution function.
+The retarded and advanced spectral functions are not actually independent.
 Otherwise, irregular divergences appear in the left-hand side of Eq.
@@
 $f^{\rm R}_{j\epsilon\pm}=-f^{\rm A}_{j-\epsilon\mp},\quad\theta^{\rm Rq}_{j\epsilon\pm}=-\theta^{\rm Aq}_{j-\epsilon\mp}.$ (13)
-Using Eqs. (11)--(14) and performing analytic continuation, we obtain the modified Usadel equations at the first and third orders of $\Delta_{\rm c}$ as follows:
+Using Eqs. (10)-(13) and performing analytic continuation, we obtain the modified Usadel equation [30] in the first and third orders of $\Delta_{\rm cl}$ as follows:
 $\mathcal{S}_{\mathrm{kernel},j}[f_{1n\pm},f_{3n\pm}]=\mathcal{S}_{\mathrm{source},j}[f_{1n\pm},f_{3n\pm}],$ (14)
 where $f_{jn\pm}=if^{\rm R}_{ji\omega_{n}\pm}$, $\omega_{n\pm}=\omega_{n}\pm ih$, $\omega_{n}=(2n+1)\pi T~(n\in\mathbb{Z})$, and $\mathcal{S}_{\mathrm{source},1}=1$.
-The others are shown in Supplemental Material [35] to avoid describing the lengthy equations, and we find that the Cooper interaction terms are not involved in the equation determining $f_{1n\pm}$.
+The others are shown in Supplemental Material [42] to avoid describing lengthy equations.
+We find that the correction from the long-range Coulomb interaction is much greater than that from the Cooper interaction.
 The equations are solved numerically by the Padé decomposition method [43], which can achieve fast convergence when calculating the Matsubara summation.
-Equation of state: The gap equation is obtained from the stationary condition $[\partial S/\partial\Delta_{\rm q}]_{\Delta_{\rm q}\rightarrow 0}=0$.
-In GL theory, the equation has the form $\alpha_{2}\Delta_{\rm c}+\alpha_{4}\Delta_{\rm c}^{3}=0$, where
-$\alpha_{2}$ $=4\,\Gamma^{-1}_{\rm c}\!-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{[}f_{1n\lambda}\!+\!\frac{1}{\omega_{n}\!+\!i\lambda h}\Big{]}$
+Equation of state: The gap equation is obtained from the stationary condition $\partial S/\partial\Delta_{\rm q}]_{\Delta_{\rm q}\rightarrow 0}=0$.
+At the GL level, the equation is in the form $\alpha_{2}\Delta_{\rm cl}+\alpha_{4}\Delta_{\rm cl}^{3}=0$, which corresponds to minimizing the GL free energy $f_{\mathrm{GL}}=\mathcal{C}\,(\frac{1}{2}\alpha_{2}\Delta^{2}_{\rm cl}+\frac{1}{4}\alpha_{4}\Delta^{4}_{\rm cl})$, where $\mathcal{C}$ is a numerical constant.
+The coefficients $\alpha_{2}$ and $\alpha_{4}$ are given as follows:
+$\alpha_{2}$ $=4\,\Gamma^{-1}_{\rm c}\!-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big[f_{1n\lambda}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{1}{\omega_{n}\!+\!i\lambda h}\Big]$
 $\quad-2\int_{0}^{\omega_{\rm D}}\frac{d\epsilon}{\epsilon}\Big[\tanh\left({\frac{\epsilon_{+}}{2T}}\right)+\tanh\left({\frac{\epsilon_{-}}{2T}}\right)\Big],$ (15)
-$\alpha_{4}$ $=-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{(}f_{3n\lambda}-\!\frac{1}{6}[f_{1n\lambda}]^{3}\Big{)}.$ (17)
+$\alpha_{4}$ $=-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big(f_{3n\lambda}-\!\frac{1}{6}[f_{1n\lambda}]^{3}\Big),$ (16)
+where $\Gamma_{\rm c}$ is the SC coupling constant.
 Note that $f_{jn\lambda}$ is derived from Eq.
 (14), and $\alpha_{2}$ is regularized by introducing the ultraviolet cutoff, which is the Debye energy $\omega_{\rm D}$, to the energy integration.
-The form of the gap equation in the GL level suggests that the free energy and the supercurrent are given as $f_{\mathrm{GL}}=\frac{1}{2}\alpha_{2}\Delta^{2}_{\rm c}+\frac{1}{4}\alpha_{4}\Delta^{4}_{\rm c}$ and $I_{\rm s}=\partial f_{\mathrm{GL}}/\partial A_{y}$, respectively.
-The helical SC state for any supercurrent $I_{s}$ is then determined by the latter equation.
-In particular, the second-order SC phase transition line in the absence of the supercurrent is calculated based on the conditions $\alpha_{2}=0$ and $\partial\alpha_{2}/\partial A_{y}=0$.
-From now on, let $A_{0}$ be the $A_{y}$ field in the zero-current SC states.
+Since we are considering helical SC states, the supercurrent $I_{s}$ in the $y$-direction is given as follows [4]:
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{\rm s}=\partial f_{\mathrm{GL}}/\partial A_{y}=-\frac{1}{4}\mathcal{C}\frac{\partial}{\partial A_{y}}\,(\alpha^{2}_{2}/\alpha_{4}).$} (17)
+The second-order SC phase transition line in the absence of the supercurrent is calculated based on the conditions $\alpha_{2}=0$ and $\partial\alpha_{2}/\partial A_{y}=0$.
 We denote the transition temperature at a magnetic field $h$ as $T_{\rm c}(h)$, which is denoted as $T_{\rm c0}(h)$ when we neglect the e-e interaction.
-At $h=0$, we describe $T_{\rm c0}(0)=T_{\rm c0}$ and $T_{\rm c}(0)=T_{\rm c}$, for simplicity.
-For later use, we define the reduced temperature $\tau_{\rm r}$ as $\tau_{\rm r}=T/T_{\rm c0}$ or $\tau_{\rm r}=T/T_{\rm c}$.
-WL conductivity: When the current reaches its critical value, the SC state changes to a resistive state.
-Here, we study the WL correction to the conductivity, $\sigma_{\rm WL}$ [37, 38, 39].
-For simplicity, we neglect the e-e interactions.
-We find that it is convenient to work with the alternative normal saddle point $\bar{Q}^{\prime}_{\rm N}=U^{\prime}\Lambda^{\prime}_{\rm N}U^{\prime-1}$, where $\Lambda^{\prime}_{\rm N}=\mathrm{diag}\,\{\gamma_{3},\gamma_{3}\}$ and $U^{\prime}=\mathrm{diag}\,\{\check{U}_{\rm K},\check{U}^{\mathbb{T}}_{\rm K}\}$.
-Here, $U^{\mathbb{T}}$ denotes the transpose of a matrix and the interchange of the time of $U$, which means that $U^{\mathbb{T}}_{t_{1}t_{2}}=U^{\rm t}_{t_{2}t_{1}}$ with $U^{\rm t}_{t_{2}t_{1}}$ being the ordinary transpose in matrix elements of $U_{t_{2}t_{1}}$.
-The matrix $\check{U}_{\rm K}$ has the same form as Eq.
-(4), but $\tau_{0}$ is replaced by $\sigma_{0}$.
-This saddle point comes from the time-reversal symmetry condition of the matrix field $Q$.
-To the best of our knowledge, this saddle point was first introduced in Ref.
-40 to calculate the WL conductivity $\sigma_{\rm WL}$ in the spinless case within the Keldysh functional formalism.
-Here, we generalize their formalism to include spin degrees of freedom.
-We explain the procedure in [35], and the equation for $\sigma_{\rm WL}$ is consistent with existing results [41, 42, 44, 43].
-Numerical results: In the numerical calculations, we fix the parameters, $\Gamma_{\rm c}=0.2$, $\omega_{\rm D}/T_{\rm c0}=0.1309$, and Fermi energy $E_{\rm F}/T_{\rm c0}=p_{\rm F}^{2}/(2mT_{\rm c0})=10^{3}$, and consider the following two sets of parameters: (I) $\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$ and (II) $\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$, where $\alpha$ is the Rashba SOC constant and $\tau$ is the impurity scattering time.
-Note that the values of $\tau T_{\rm c0}$ in cases (I) and (II) correspond to the characteristic lengths of $l_{\rm F}p_{\rm F}=30$ and 151515, respectively, which are consistent with the WL regime.
-The characteristic energies of SOC and disorder are represented by $\epsilon_{\rm SOC}=\alpha v_{\rm F}$ and $\epsilon_{\rm D}=1/\tau$, respectively.
-In case (I), the SOC is comparable with disorder as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.6$, while case (II) is in a weak SOC regime as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.06$.
-We show the numerical results for the comparable and weak SOC in Figs.
-1 and 2, respectively.
-Figure 1: The diode quality factor $\eta$ and the WL correction to conductivity $\sigma_{\rm WL}$ in the intermediate SOC regime ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$).
-The solid (dashed) lines show the results in the interacting (free) case with (without) the e-e interaction.
-Panels (a) and (c) show the color maps of $\sigma_{\rm WL}$ and $\eta$, respectively.
-The solid and dashed lines show the second-order superconducting transition line.
-The bold circles in (a) represent the tricritical points.
-Panels (b) and (d) show $\sigma_{\rm WL}$ and $\eta$ near the SC transition line, respectively.
-In these figures, we vary the reduced temperature $\tau_{\rm r}$ and magnetic field $h$ so that the temperature is $T=0.99T_{\rm c}(h)$.
-We first show the properties of zero-current SC states.
+For simplicity we write $T_{\rm c0}(0)=T_{\rm c0}$ and $T_{\rm c}(0)=T_{\rm c}$.
+Furthermore, we calculate the critical currents $I_{\rm sc\pm}$ in the $\pm y$-direction, which is obtained by optimizing $I_{\rm s}$, with focus on the diode quality factor defined as $\eta=(1-\xi)/(1+\xi)$ with $\xi=|I_{\rm sc-}/I_{\rm sc+}|$.
+In the numerical calculations, we fix the parameters, $\Gamma_{\rm c}=0.2$, $\omega_{\rm D}/E_{\rm F}=0.1309$, and Fermi energy $E_{\rm F}/T_{\rm c0}=p_{\rm F}^{2}/(2mT_{\rm c0})=10^{3}$, where $p_{\rm F}=mv_{\rm F}$ and $m$ is the electron mass.
+For the Rashba SOC constant $\alpha$ and impurity scattering time $\tau$, we consider the following two sets of parameters: (I) $\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$ and (II) $\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$.
+Note that the characteristic energies of SOC and disorder are represented by $\epsilon_{\rm SOC}=\alpha v_{\rm F}$ and $\epsilon_{\rm D}=1/\tau$, respectively.
+Therefore, case (I) is considered a comparable SOC case with $\epsilon_{\rm SOC}/\epsilon_{\rm D}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.3}$, while case (II) is considered a weak SOC case with $\epsilon_{\rm SOC}/\epsilon_{\rm D}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.03}$.
+Note also that our perturbative approach is valid for large dimensionless conductance $g\gg 1$ [38, 30].
+Here, $g$ is given as $g=2\pi\sigma_{\rm D}$ [40], where $\sigma_{\rm D}=2\nu D$ is the Drude conductivity and $\nu=m/(2\pi)$ is the 2D density of states per spin.
+We have confirmed that the impurity scattering time $\tau$ in both parameter sets (I) and (II) satisfies this condition.
+Figure 1: The phase diagram and diode quality factor $\eta$ in the comparable SOC case ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$).
+The solid (dashed) lines show the results in the case with (without) e-e interactions, which is referred to as the interaction (free) case in the main text.
+(a) The color maps of $\eta$ in the vicinity of the second-order SC transition lines, which change to first-order transitions at the tricritical points represented by the black circles.
+(c) The transition lines from panel (a) on the normalized temperature-normalized magnetic field scale, namely the $\tau_{\rm r}$-$h_{\rm r}$ scale.
+(b) and (d) the diode quality factor $\eta$ as a function of $\tau_{\rm r}$ and $h_{\rm r}$, respectively.
+In these figures, we vary $\tau_{\rm r}$ and $h_{\rm r}$ so that the temperatures satisfy $T=0.99T_{\rm c0}(h)$ in the free case and $T=0.99T_{\rm c}(h)$ in the interaction case.
+Note that the same $\tau_{\rm r}$ does not imply the same $h$ or $h_{\rm r}$ between the interaction and free cases.
+Phase diagram: We first show the properties of zero-current SC states.
 As shown in Figs.
-1a and 2a, the e-e interactions suppress the transition temperatures and critical magnetic fields.
-The gauge field $A_{0}$ corresponding to the momentum of Cooper pairs is also suppressed when the interactions are incorporated [35].
-Since there is no contribution from the Cooper channel in the equation determining the second-order SC transition line, the suppression of the transition temperature is attributed to the Coulomb channel.
+1a and 3a, the transition temperatures and magnetic fields are lower in the interaction case than in the free case.
+The gauge field of the helical SC state at zero current is also suppressed when the interactions are incorporated [42].
 The zero-field transition temperatures are $T_{\rm c}=0.8793T_{\rm c0}$ and $T_{\rm c}=0.6457T_{\rm c0}$ in Figs.
-1 and 2, respectively.
+1a and 3a, respectively.
 We have confirmed that $T_{\rm c}$ decreases as the impurity scattering time $\tau$ decreases, while it increases toward $T_{\rm c0}$ as the SOC $\alpha$ increases.
-The transition lines depicted in Figs.
-1a and 2a show the second-order phase transition, and they terminate at the tricritical points at high magnetic fields, indicating the first-order transition, at which $\alpha_{4}$ becomes negative.
-It is well known that the first-order SC transition occurs in the high field region due to the Pauli paramagnetic effects when we adopt the mean-field theory for spatially uniform $s$-wave states [45].
-At magnetic fields higher than the tricritical point, the second-order transition from the normal state to the non-uniform stripe SC state is expected [46].
-Since the Cooper channel makes a non-zero contribution to $\alpha_{4}$, both Coulomb and Cooper channels affect the tricritical points.
-Figure 2: The WL correction to conductivity $\sigma_{\rm WL}$ in the weak SOC regime ($\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$).
-Panel (a) shows the color map of $\sigma_{\rm WL}$, which is plotted in the same manner as Fig.
+The second-order transition lines become first-order transitions at the tricritical points, where $\alpha_{4}$ becomes negative.
+The tricritical temperature and magnetic field in the interaction case, $T^{*}_{\rm c}$ and $h^{*}_{\rm c}$, are lower than those in the free case, $T^{*}_{\rm c0}$ and $h^{*}_{\rm c0}$, respectively.
+This first-order transition is well known to occur in the high field region due to the Pauli paramagnetic effects [52], and it also occurs when the SOC is sufficiently weak, namely $\Gamma_{\rm r}=2D\alpha^{2}\lesssim 10\Delta_{0}$ where $\Delta_{0}\sim 1.8T_{\rm c0}$ is the SC gap at zero temperature [22, 53].
+The parameters used here satisfy this condition.
+At magnetic fields higher than the tricritical points, although a second-order transition to a stripe SC state is expected in the clean limit [54], it turns out to be a first-order transition to a uniform SC state in the dirty limit [55].
+Unlike the previous study [22] that solves the MF gap equation rigorously, our approach cannot describe the first-order transition because the $\Delta_{\rm cl}$-term in the free energy is up to the 4th order.
+Figure 2: The diode quality $\eta$ is plotted as a function of the normalized magnetic field $h_{\rm r}$ at normalized temperatures: (a) $\tau_{\rm r}=0.9$ and (b) $\tau_{\rm r}=0.25$.
+In panel (a), the red and blue squares denote $\eta$ in the interaction and free cases, respectively.
+The green line shows the fitting by Eq.
+(18).
+In panel (b), the red solid and blue dashed lines show $\eta$ in the interaction and free cases, respectively.
+SD effect: Let us discuss the SD effect with focus on the diode quality factor $\eta$.
+Since $\eta$ is tiny in the weak SOC case, we calculate the SD effect in the comparable SOC case and show the color map of $\eta$ in Fig.
 1a.
-Panel (b) shows $\sigma_{\rm WL}$ near the superconducting transition line in the same manner as Fig.
-1b.
-Next, let us focus on the quality factor $\eta$ of the SD effect.
-The diode quality factor $\eta$ is defined as $\eta=(1-\xi)/(1+\xi)$, where $\xi=|I_{\rm sc-}/I_{\rm sc+}|$ and $I_{\rm sc\pm}$ is the critical current in the ±$\pm y$-direction.
-Because $\eta$ is tiny in the weak SOC regime, we calculate the SDE in the intermediate SOC regime and show the results of $\eta$ in Figs.
-1c and 1d.
-As shown in Fig.
-1c, the diode quality factor increases as the system approaches the tricritical point.
-To show the behaviors of the SDE near the transition line, we plot in Fig.
-1d the diode quality factor near the transition line as a function of the reduced temperature.
-The absolute values of the diode quality factor increase divergently as the tricritical point is approached.
-We see qualitatively the same behaviors in the absence and presence of the e-e interaction, while the tricritical point is shifted.
-Although the Coulomb interaction significantly reduces the zero-field transition temperature, the diode quality factor almost coincides in both cases at high reduced temperatures.
-This behavior is deviated as the tricritical points are approached.
-Because the tricritical point is shifted by interactions, the absolute values of the diode quality factor in the presence of interactions become smaller than those in the case without interactions.
-Finally, we discuss the WL correction to the conductivity $\sigma_{\rm WL}$.
-The correction is normalized with the Drude conductivity $\sigma_{\rm D}=2\nu D$.
-As shown in Fig.
-1a, interestingly, $\sigma_{\rm WL}$ in the intermediate SOC regime decreases as the temperature and magnetic field increase, indicating stronger localization effects.
-By examining the plot near the transition lines [Fig.
-1b], we conclude that the e-e interaction can suppress the WL correction.
-However, as shown in Fig.
-2a, the qualitatively opposite behaviors manifest themselves in almost the entire region of temperatures and magnetic fields.
-Looking at the results near the transition lines [Fig.
-2b], we find that the e-e interaction can enhance the WL correction when the SOC is weak.
-Summary: We have formulated the renormalized GL theory including the Cooper and Coulomb interaction terms from the NLSM based on the Keldysh functional formalism.
+It appears that $\eta$ increases as the system approaches the tricritical point.
+To show the behavior of $\eta$ near the transition line, we plot in Fig.
+1b the diode quality factor near the transition line as a function of normalized temperature $\tau_{\rm r}$, defined as $\tau_{\rm r}=T/T_{\rm c0}$ or $\tau_{\rm r}=T/T_{\rm c}$.
+We see qualitatively the same behavior of $\eta$ in the absence and presence of the e-e interaction.
+Especially, $\eta$ in the two cases almost coincide at high $\tau_{\rm r}$ despite the significant decrease in the zero-field transition temperature due to the e-e interactions.
+This means that the SD effect with the same order of $\eta$ as in the free case can be achieved in the interaction case.
+This finding also suggests a universal law of the diode quality factor, which will be derived below.
+In the GL theory [3, 4], the diode quality factor near the second-order transition line follows the scaling law $\eta\sim\mathcal{F}\left[{1-\tilde{\tau}}\right]^{1/2}$.
+In the free case, the reduced temperature is $\tilde{\tau}=T/T_{\rm c0}(h)$, and the coefficient can be expanded by odd powers of the magnetic field, $\mathcal{F}=\sum_{n=0}A_{2n+1}h^{2n+1}$.
+In the interaction case, we replace $T_{\rm c0}(h)$ and $A_{2n+1}$ with $T_{\rm c}(h)$ and $A^{\prime}_{2n+1}$, respectively.
+The relation between $A_{1}$ and $A^{\prime}_{1}$ is derived from the following analysis.
+We can fit transition lines in Fig.
+1a by $T_{\rm c0}(h)/T_{\rm c0}=[1-\chi_{\rm c0}(h/h_{\rm c0})^{2}]^{1/\chi_{\rm c0}}$ and $T_{\rm c}(h)/T_{\rm c}=[1-\chi_{\rm c}(h/h_{\rm c})^{2}]^{1/\chi_{\rm c}}$.
+These reduce to the known behavior of the transition line $\tau_{\rm r}\simeq 1-h^{2}_{\rm r}$ at low fields, where $h_{\rm r}\equiv h/h_{\rm c0}$ $(h/h_{\rm c})$ and $\tau_{\rm r}=T/T_{\rm c0}$ $(T/T_{\rm c})$.
+Thus, the transition lines almost coincide on the $\tau_{\rm r}$-$h_{\rm r}$ scale at low fields [Fig.
+1c].
+We find $h_{\rm c0}=3.61T_{\rm c0}$, $h_{c}=3.28T_{\rm c0}$, $\chi_{\rm c0}=1.89$, and $\chi_{c}=1.84$.
+Note that $h_{\rm c0}$ and $h_{\rm c}$ are not equivalent to the tricritcal fields ($h^{*}_{\rm c0}=2.6T_{\rm c0}$ and $h^{*}_{\rm c}=2.4T_{\rm c0}$).
+When the horizontal axis in Fig.
+1b is changed to $h_{\rm r}$, the diode quality factors nearly overlap at low fields, as shown in Fig.
+1d.
+This means that $A_{1}/h_{\rm c0}=A^{\prime}_{1}/h_{\rm c}\equiv\mathcal{A}_{1}$, and $\eta$ at low fields follows the universal behavior
+$\eta=\mathcal{F}\big[1-\tilde{\tau}\big]^{1/2},\quad\mathcal{F}\simeq\mathcal{A}_{1}h_{\mathrm{r}}.$ (18)
+We further find that this law is valid in a wide range of the phase diagram at high temperatures $\tau_{\rm r}>0.9$, as demonstrated in Fig.
+2a by setting $\mathcal{A}=2.8\times 10^{-3}\tau^{-1}_{\rm r}$.
+Based on these analyses, we conclude that the diode quality factor is robust against the corrections due to disorder and e-e interactions in the low-magnetic-field regime.
+In previous work [12], a system essentially similar to ours was studied without incorporating e-e interactions.
+The linear field dependence of $\eta$ is found to be due to the singlet-triplet mixing mechanism.
+However, unlike our Eq.
+(18), the analytical expression derived in Ref.
+[12] cannot describe the whole behavior of $\eta$ in Fig.
+2a.
+Nevertheless, combining their results and Eq.
+(18) suggests that the interactions may lower both Rashba SOC and transition temperature while preserving the SD effect.
+In the low-temperature regime, deviation from the universal law manifests, as shown in Fig.
+2b.
+The roles of the e-e interactions appear non-monotonically.
+In the high-magnetic-field regime, the diode quality factor $\eta$ is higher in the interaction case than in the free case.
+This is a result of an increase in the tricritical field in the $h_{\rm r}$ scale due to e-e interactions, and can be qualitatively explained by Eq.
+(18).
+However, at low magnetic fields, the diode quality factor is decreased by the interaction.
+WL conductivity: When the supercurrent reaches its critical value, the SC state changes to a resistive state.
+Here, we study the quantum corrections to conductivity in the resistive states.
+For simplicity, we neglect the corrections to conductivity due to e-e interactions and consider the conventional WL conductivity $\sigma_{\rm WL}$ [44, 45, 46].
+The calculation of $\sigma_{\rm WL}$ can be performed in the Keldysh functional formalism by extending the previous work [47] that considers only the spinless case to include the spin degrees of freedom [42].
+It appears that our equation for $\sigma_{\rm WL}$ is consistent with existing results [48, 49, 51, 50].
+The numerical results of the WL conductivity $\sigma_{\rm WL}$ are shown in Fig.
+3.
+The WL correction normalized by the Drude conductivity is much larger in the weak SOC case [Fig.
+3a] than in the comparable SOC case [Fig.
+3b].
+As the temperature and magnetic field increase, the magnitude of $\sigma_{\rm WL}$ decreases in the weak SOC case but increases in the comparable SOC case.
+These two results suggest that the localization behaviors are suppressed by the SOC irrespective of Coulomb interactions.
+In contrast, the magnitude of the diode quality factor $\eta$ with a large SOC is usually larger than with a small SOC for a given temperature and magnetic field.
+Therefore, our results reveal a trade-off relationship between the SD effect and the WL behavior in the resistive states, at least in the parameter range considered here.
+Figure 3: The WL conductivity $\sigma_{\rm WL}$ in (a) the weak SOC regime ($\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$) and (b) the comparable SOC case ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$).
+The solid (dashed) lines show the results in the interaction (free) case.
+The white circles represent the tricritical points.
+Summary: We have formulated the renormalized GL theory including the Cooper and long-range Coulomb interaction terms from the NLSM, based on the Keldysh functional formalism.
 Our formalism extends the previous one and enables us to study dirty 2D Rashba superconductors under in-plane Zeeman fields, which are typical platforms of the SD effect.
-We find that the interplay between localization and interactions leads to the corrections that suppress the SC states, the first-order transition, and the SD effect.
-Furthermore, we have calculated the WL correction to conductivity that characterizes transport when the SC state is destroyed by an electric current exceeding the critical current.
-The WL correction can be suppressed or enhanced by interactions depending on the strength of SOC.
-We expect our results to pave the way for switching between superconducting, metallic, and insulating transport via an electric current.
+We found the robustness of the SD effect with respect to the Cooper and Coulomb interactions near the low-transition magnetic fields, even though the SC states and first-order transition are suppressed by the long-range Coulomb interaction in particular.
+This means that the magnetic field required to produce a diode quality factor can be reduced by the Coulomb interaction.
+We also demonstrated a trade-off relationship between the localization behavior in the resistive state, which emerges above the critical current, and the efficiency of the SD effect.
+We expect our results to pave the way for switching between superconducting, metallic, and insulating behavior via an electric current.
 Acknowledgements — We appreciate helpful discussions with Ryusuke Ikeda, Hideki Narita, Teruo Ono, and Ryuji Hakuno.